OpenStudy (anonymous):

if f(x) = 2x^2 and r(x) = 2x^(1/2), find a simplified formula for f(r(x))implified

OpenStudy (anonymous):

Imput the equation of r(x) into the f(x) equation and simplify. f(r(x))= 2(2x^1/2)^2 since you multiply exponents when an exponent is raised, and 1/2x2=1, the exponents cancel. (any number raised to the first power is that number) but don't forget to square the 2 first! f(r(x))=2(2^2x) f(r(x))=2(4x) f(r(x))=8x I hope that helps!